Dewar temperatures and bias voltages/currents

march,2002

 dewar temperatures
dewar bias currents and voltages

Dewar temperatures:

   The receiver systems have temperature sensors that measure the physical temperature of the 16 Kelvin, 70 Kelvin, and the OMT stages of the receiver. These temperatures are recorded daily. They are plotted below by year and for the last 5 and 30 days (updated at 5:25 am each day). The color plots overplot all of the receivers on one plot using color to differentiate between the receviers. The black and white plots have a mosaic with 1 plot per receiver (if you don't have a color plotter, then you can make hardcopy with the black and white version). Some receivers do not have the temperature diodes to measure the 70K or the OMT stages. These plots will appear blank. See using idl to plot the dewar temperatures for routines that allow you to plot this data interactively.
 
plots of dewar temperature verus time
Date
16K stage
color
70K stage
color
OMT stage
color
16K stage
black & white
70K stage
black & white
OMT stage
black & white
last 5 days
16k_last5
70K_last5
omt_last5
16K_last5
70K_last5
omt_last5
last 30 days
 16K_last30
70K_last30
omt_last30
16K_last30 
70K_last30
omt_last30
Year to date
16K_2018
70K_2018
omt_2018
16K_2018
70K_2018
omt_2018
2017
16K_2017
70K_2017
omt_2017
16K_2017
70K_2017
omt_2017
2016
16K_2016
70K_2016
omt_2016
16K_2016
70K_2016
omt_2016
2015
16K_2015
70K_2015
omt_2015
16K_2015
70K_2015
omt_2015
2014
16K_2014
70K_2014
omt_2014
16K_2014
70K_2014
omt_2014
2013
16K_2013
70K_2013
omt_2013
16K_2013
70K_2013
omt_2013
2012
16K_2012
70K_2012
omt_2012
16K_2012
70K_2012
omt_2012
2011
16K_2011
70K_2011
omt_2011
16K_2011
70K_2011
omt_2011
2010
16K_2010
70K_2010
omt_2010
16K_2010
70K_2010
omt_2010
2009
16K_2009
70K_2009
omt_2009
16K_2009
70K_2009
omt_2009
2008
16K_2008
70K_2008
omt_2008
16K_2008
70K_2008
omt_2008
2007
16K_2007
70K_2007
omt_2007
16K_2007
70K_2007
omt_2007
2006
16K_2006
70K_2006
omt_2006
16K_2006
70K_2006
omt_2006
2005
16K_2005
70K_2005
omt_2005
16K_2005
70K_2005
omt_2005
2004
16K_2004
70K_2004
omt_2004
16K_2004
70K_2004
omt_2004
2003
16K_2003
70K_2003
omt_2003
16K_2003
70K_2003
omt_2003
2002
16K_2002
70K_2002
omt_2002
16K_2002
70K_2002
omt_2002
2001
16K_2001
70K_2001
omt_2001
16K_2001
70K_2001
omt_2001
2000
16K_2000
70K_2000
omt_2000
16K_2000
70K_2000
omt_2000
1999
16K_1999
70K_1999
omt_1999
16K_1999
70K_1999
omt_1999


Dewar bias voltages and currents

    The dewar monitoring system includes monitoring of the bias voltages and currents for each stage of the amplifiers. The data is recorded (all receivers are read about once every 30 seconds) and written to discs. Plots are made daily  (at 5:25am) of: the last 5 days, the last 30 days, and the year to date.
The plots contain:


Bias Currents and voltages for each receiver
Date
327
430
800
lbw
sbn
sbw
sbh
cb
cbw
cbh
xb
last 5 days
.pdf .pdf
.pdf .pdf .pdf .pdf .pdf

.pdf
last 30 days
.pdf .pdf
.pdf .pdf .pdf .pdf .pdf

.pdf
Year to date
.pdf .pdf
.pdf .pdf .pdf .pdf .pdf .pdf
.pdf
2017
.pdf .pdf
.pdf .pdf .pdf .pdf .pdf .pdf .pdf .pdf
2016
.pdf .pdf
.pdf .pdf .pdf .pdf .pdf
.pdf .pdf
2015
.pdf .pdf
.pdf .pdf .pdf .pdf .pdf
.pdf .pdf
2014
.pdf .pdf
.pdf .pdf .pdf .pdf .pdf
.pdf .pdf
2013 .pdf .pdf
.pdf .pdf .pdf .pdf .pdf
.pdf .pdf


Notes:

2013

2008

  • 08mar04 - cbh now included in plots.
  • jun09      - cbh dropped, 800Mhz receiver monitored
  • 08aug09 - sbh dropped cbh now monitored
  • 2003

  • 13jan03- lbn dewar warmed up for maintenance. xband 16K stage drops by 5K since now 2 rather than 3 dewars on the compressor.
  • 15jan03- lbn op-amp for tempmonitor conversion I to V added bypass capacitors. Drifting should settle down..
  • 16jan03- xb,sbn,lbn 1 compressor changed to (xb,sbh)  (sbn,lbn)
  • 26mar03   - xb,sbn,lbn back to 1 compresso sbh on 1 by itself.
  • 04apr03 - lbn changed diode monitoring circuit to use old diode curve.
  • 07apr03  - changed sbn diode board to use old diode curve..updated software

  • 2002

  • 18feb02 xb - xband moved to share compressor with sbn,lbn. 16K stage jumped a few degrees.
  • 06apr02 lbw- left vacuum pump running on dewar. 70K 120->80K. must be a leak in dewar.
  • 27jun02 lbw - warmed up lbw, regreased seals, added vacion pump (but not on yet), then cooled down.
  • 10jul02 lbw  - turned on vacion pump.
  • 17sep02 lbn,xb - sbn no longer on compressor with xb,lbn (down for repairs).  xb,lbn temps decreased.
  • 24sep02 lbn,xb,sbn back to being on same compressor.
  • 20dec02 switched to use continous sampling. last5 plots sampled every 22 secs, last_30 plots smoothed to 4 minutes.
  • processing: x101/dwtemp/dwtemp.pro
     home_~phil